# Enhancing capabilities of particle image/tracking velocimetry with physics-informed neural networks

### v2

@@ Abstract
Solving inverse problems, which means obtaining model parameters from observed data, using conventional computational fluid dynamics solvers is prohibitively expensive. Here we employ machine learning algorithms to overcome the challenge. As an example, we consider a moderately turbulent fluid flow, excited by a stationary force and described by a two-dimensional Navier-Stokes equation with linear bottom friction. Given sparse and probably noisy data for the velocity and the general form of the model, we reconstruct the dense velocity and pressure fields in the observation domain, infer the driving force, and determine the unknown fluid viscosity and friction coefficient. Our approach involves training a physics-informed neural network by minimizing the loss function, which penalizes deviations from the provided data and violations of the Navier-Stokes equation. The suggested technique extracts additional information from experimental and numerical observations, potentially enhancing the capabilities of particle image/tracking velocimetry.
@@ I Introduction
Extracting information and reconstructing a flow field from sparse or noisy measurements is important yet difficult in many applications. Solving the problem requires additional knowledge to compensate for imperfections in the data. A well-constructed physical model of the phenomenon can serve in this capacity. However, traditional computational fluid dynamics (CFD) models are ill-suited for these purposes due to the computational complexity of data assimilation methods Asch, Bocquet, and Nodet (2016), especially for turbulent flows. Recent advances in machine learning (ML) open up new opportunities for addressing these challenges Brunton, Noack, and Koumoutsakos (2020); Cai et al. (2021a); Karniadakis et al. (2021); Sharma et al. (2023).
In this work, we focus on physics-informed neural networks (PINNs) Lagaris, Likas, and Fotiadis (1998); Raissi, Perdikaris, and Karniadakis (2019), which provide a framework for merging data with physical laws expressed as partial differential equations (PDEs). Previously, PINNs were applied to uncover near-wall blood flow from sparse data Arzani, Wang, and D’Souza (2021), for super-resolution and denoising applications Fathi et al. (2020); Gao, Sun, and Wang (2021); Eivazi and Vinuesa (2022), to reconstruct velocity and pressure fields from flow visualizations Raissi, Yazdani, and Karniadakis (2020); Cai et al. (2021b); Clark Di Leoni et al. (2023a) and particle image/tracking velocimetry (PIV/PTV) data Wang, Liu, and Wang (2022); Clark Di Leoni et al. (2023b), to simulate turbulence Jin et al. (2021); Kag, Seshasayanan, and Gopinath (2022); Eivazi et al. (2022), just to name a few. A more comprehensive bibliography can be found in recent reviews Cai et al. (2021a); Karniadakis et al. (2021); Cuomo et al. (2022); Sharma et al. (2023). Here we implement PINN to solve inverse problem inspired by experimental studies of forced turbulent flows in thin conducting fluid layers Sommeria (1986); Paret and Tabeling (1998); Boffetta et al. (2005); Xia, Shats, and Falkovich (2009); Orlov, Brazhnikov, and Levchenko (2018); Fang and Ouellette (2021). Our goal is to show that when the physical model is taken into account, the suggested technique has the ability to extract more information from conventional flow observations.
@@
We consider two-dimensional incompressible fluid flow described by the forced Navier-Stokes equation with linear bottom friction
@@
where $\bm{v}$ is 2D velocity, $p$ is the pressure, $\alpha$ is the friction coefficient, and $\nu$ is the kinematic viscosity. The external force $\bm{f}$ is assumed to be stationary in time and divergence-free in space. We are interested in a moderately turbulent regime, when the nonlinear term in the Navier-Stokes equation plays a substantial role. Then, despite the stationary nature of the external forcing, the fluid flow is characterized by strong spatiotemporal variations. Now suppose that someone measures the velocity field in a certain region (possibly with open boundaries) and makes observations for several times longer than the typical time of flow fluctuations. In what follows, we will demonstrate how to reconstruct the dense velocity $\bm{v}(\bm{r},t)$ and pressure $p(\bm{r},t)$ fields, and establish the driving force $\bm{f}(\bm{r})$ and unknown fluid viscosity $\nu$ and bottom friction $\alpha$, based only on potentially sparse and noisy velocity field data. Let us emphasize that the developed approach can be easily implemented on top of the existing PIV/PTV workflows, thereby expanding their functionality.
The contribution of this paper compared to previous studies is twofold. First, we generate a novel dataset of forced two-dimensional fluid flow that can serve as a benchmark solution in future research Hao et al. (2023). Second, we extend the PINN algorithm by including the ability to infer driving forces. This is achieved by simultaneously training two neural networks (NNs), one for the velocity and pressure field and the other for the external force. This division, in particular, makes it possible to explicitly take into account the stationarity of the external forcing in the design of the corresponding NN. Including physical information in NN architecture is known to enhance training performance and prediction accuracy.
II Data generation
@@
Data for this study are obtained by integrating (1) using the GeophysicalFlows.jl pseudospectral code Constantinou et al. (2021), which has recently been successfully applied to model two-dimensional turbulence Parfenyev (2022); Kolokolov, Lebedev, and Parfenyev (2024); Parfenyev (2024). The code can be executed on the GPU, resulting in high computational performance. The aliasing errors are removed with the two-third rule. The simulation domain is a doubly periodic box of size $2\pi\times 2\pi$. The boundary conditions and form of the domain are only relevant to the data production process; the rest of the study is generic in this regard.
The observation area is smaller and has the size of $L\times L$ with $L=\pi$, see Fig. 1a. Thus, fluid flows in and out of the observation zone and the movement within it depends on the unseen surroundings. The system parameters $\nu=0.01$ and $\alpha=0.1$ are chosen to be consistent with the experimental studies. The external forcing has the Kolmogorov form
@@
with $f_{0}=10$ and $k_{f}=5$. We start from an arbitrary random initial condition, and after a transient process, the flow reaches a statistical steady-state. Once stationary, we saved data at $\Delta t=0.02$ intervals for $T=4$ units, covering several turn-over times of flow fluctuations. The spatial resolution is $512^{2}$ for the entire computational domain and $256^{2}$ for the observation zone.
@@
Fig. 1a shows snapshot of the vorticity field $\omega=\partial_{x}v_{y}-\partial_{y}v_{x}$ in the statistical steady-state. The flow is chaotic, with no obvious imprint of an external forcing. Since there are two dissipation mechanisms – viscosity and bottom friction – the flow state can be characterized by two dimensionless parameters Mishra et al. (2015): the Reynolds number $Re=UL/\nu\approx 1.3\times 10^{3}$ and $Rh=U/(\alpha L)\approx 13$, which were computed using the root mean square velocity $U$ and the size $L$ of the observation domain. The energy spectrum is presented in Fig. 1b. It indicates that fluid flow is mainly determined by modes with wave vectors $k\lesssim k_{f}$.
Next, we assume that measurements reveal the velocity field $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$ at $N_{data}$ points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, which are randomly scattered within the observation area and time interval of length $T$. These data imitate PIV/PTV measurements and will be used for subsequent neural network training. To study the influence of measurement noise, we will distort the velocity as $v_{x,y}^{d}\rightarrow(1+\eta)v_{x,y}^{d}$, where $\eta\sim\mathcal{N}(0,\varepsilon^{2})$ is a normally distributed random variable and the standard deviation $\varepsilon$ controls the noise level. We set $N_{data}=3\times 10^{4}$, which is around $0.2\%$ of the total data and thus represents sparse measurements. On average, this corresponds to $150$ points per snapshot compared to $65536$ points at full resolution.
III Physics-Informed Neural Network
@@
In our study, we utilize two independent NNs that are trained simultaneously, see Fig. 2a. The first NN approximates the velocity and pressure at a particular position and time. It has three input neurons and three output neurons, $(v_{x},v_{y},p)=\mathcal{F}_{\vec{\theta}_{1}}(x,y,t)$. The second NN is intended to predict stationary force. As a result, it contains only two input neurons and two output neurons, $(f_{x},f_{y})=\mathcal{G}_{\vec{\theta}_{2}}(x,y)$. The parameters $\vec{\theta}_{1}$ and $\vec{\theta}_{2}$ denote the trainable variables for NNs.
@@
Figure 3: Comparison between DNS data (ground truth) and PINN predictions for model $A_{1}$ (with clean data).
@@
In general, increasing the size of NNs can reduce the prediction error, but only when the NNs are given enough data Wang, Liu, and Wang (2022). Here we set $N_{l}=7$ and $N_{n}=250$ for $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$, and $N_{l}=5$ and $N_{n}=30$ for $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$. This choice of parameters is consistent with $N_{data}=3\times 10^{4}$. Note that $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$ is much simpler than $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$ because it is not time dependent and the external forcing usually does not have a complex spatial dependency.
@@
$\mathcal{L}_{data}=\dfrac{1}{N_{data}U^{2}}\sum_{n=1}^{N_{data}}|\bm{v}(\bm{r}_{n}^{d},t_{n}^{d})-\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})|^{2},$ (4)
@@
$\bm{e}_{1}=\left(\partial_{t}\bm{v}+(\bm{v}\nabla)\bm{v}+\nabla p+\alpha\bm{v}-\nu\nabla^{2}\bm{v}-\bm{f}\right)L/U^{2},\;$ (6)
$e_{2}=\left(\partial_{x}v_{x}+\partial_{y}v_{y}\right)L/U,$ (7)
$e_{3}=\left(\partial_{x}f_{x}+\partial_{y}f_{y}\right)L^{2}/U^{2}.$ (8)
Computing the right-hand sides involves differentiating the outputs of NNs with respect to input neurons using automatic differentiation Baydin et al. (2018). The number and location of collocation points are not limited by measurements and can be chosen arbitrarily. We set $N_{eq}=3N_{data}$, with one-third of these points corresponding to data points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, and the rest are generated randomly. Note that instead of requiring the incompressibility conditions to be satisfied, the NNs can be modified to predict the relevant stream functions Raissi, Perdikaris, and Karniadakis (2019); Kag, Seshasayanan, and Gopinath (2022). However, in this case, calculating the terms in the Navier-Stokes equation involves computation of higher-order derivatives, which significantly slows down the training process.
@@
Minimization of the loss function (9) is performed by a gradient-based optimizer that adjusts network and model parameters. As the bottom friction and fluid viscosity are positive, it is convenient to represent them in the form $\alpha=e^{\lambda_{1}}$ and $\nu=e^{\lambda_{2}}$. So, the training variables are $\{\vec{\theta}_{1},\vec{\theta}_{2},\lambda_{1},\lambda_{2}\}$. First, we train the model for $10^{5}$ epochs using Adam optimizer Kingma and Ba (2014) with learning rate of $\mu=10^{-3}$. To reduce computation time, we have implemented a mixed precision technique Micikevicius et al. (2017). Then we use L-BFGS-B optimizer Byrd et al. (1995) until the tolerance reaches machine precision. The training is carried out on a single NVIDIA A100 GPU. The code is written using the PyTorch library Paszke et al. (2019) and is openly available at [rep, ].
Figure 4: Velocity (left), pressure (middle) and force (right) probability density functions for DNS data and PINN models trained with various noise levels in the input data. All figures share the same legend. The dashed line corresponds to expression (11).
IV Results

### v3

# Inferring parameters and reconstruction of two-dimensional turbulent flows with physics-informed neural networks

@@ Abstract
Obtaining system parameters and reconstructing the full flow state from limited velocity observations using conventional fluid dynamics solvers can be prohibitively expensive. Here we employ machine learning algorithms to overcome the challenge. As an example, we consider a moderately turbulent fluid flow, excited by a stationary force and described by a two-dimensional Navier-Stokes equation with linear bottom friction. Using dense in time, spatially sparse and probably noisy velocity data, we reconstruct the spatially dense velocity field, infer the pressure and driving force up to a harmonic function and its gradient, respectively, and determine the unknown fluid viscosity and friction coefficient. Both the root-mean-square errors of the reconstructions and their energy spectra are addressed. We study the dependence of these metrics on the degree of sparsity and noise in the velocity measurements. Our approach involves training a physics-informed neural network by minimizing the loss function, which penalizes deviations from the provided data and violations of the governing equations. The suggested technique extracts additional information from velocity measurements, potentially enhancing the capabilities of particle image/tracking velocimetry.
@@ I Introduction
Experimental studies of fluid flows involve measuring their characteristics. The collected data may have low spatio-temporal resolution or be corrupted by measurement noise. Sometimes, the quantity of interest cannot be measured directly, and it has to be reconstructed from other measurements (e.g., pressure can be determined from a known velocity field Van der Kindere et al. (2019)). Various adjoint- and ensemble-variational data assimilation methods address these challenges by combining experimental data with the numerical simulations Zaki and Wang (2021); Asch, Bocquet, and Nodet (2016). The basic idea is to find the flow state that satisfies the governing Navier-Stokes equation and minimizes the deviation from observations. The fundamental downside of such approaches is the computational complexity required to run multiple numerical simulations with various initial conditions and unknown system parameters that are adjusted during optimization. Once they are found, the related flow evolution gives access to the complete system state and at much higher resolution than the original data.
Recently, an alternative approach has been proposed in which the numerical solver is replaced by a physics-informed neural network (PINN) Raissi, Perdikaris, and Karniadakis (2019). The flow estimation is also treated as a minimization problem. The neural network receives coordinates and a moment of time as input and returns flow variables as output. The loss function penalizes deviations from the available data and violations of the governing equations. Let us emphasize that the governing equations are no longer strictly obeyed; rather, their residuals are used as a penalty in the optimization problem. The main advantages of these methods are the relative simplicity of implementation and better suitability for ill-defined problems. In particular, PINNs may be used in domains with unknown boundary conditions Arzani, Wang, and D’Souza (2021), whereas numerical simulations without that knowledge are impossible.
A side-by-side comparison of both methodologies was carried out in Ref. [Du, Wang, and Zaki, 2023], where it was shown that PINNs are generally less accurate for sparse data. However, the situation may change in the future, since PINNs are a relatively new technology that is currently being actively improved. Examples of the application of this method to increasingly complex flow configurations can be found in recent reviews Cai et al. (2021a); Karniadakis et al. (2021); Sharma et al. (2023).
In this paper, we apply PINNs to a specific example of two-dimensional turbulence, which is motivated by experimental studies Boffetta et al. (2005); Xia, Shats, and Falkovich (2009); Orlov, Brazhnikov, and Levchenko (2018); Fang and Ouellette (2021). We assume that flow measurements reveal the velocity at some locations and time moments. These data can, for example, be obtained by particle image/tracking velocimetry (PIV/PTV) techniques, which track tracing particles over consecutive time frames to determine velocity. The collected data are assumed to be dense in time, but may be sparse in space and contain measurement errors. Density in time means that PIV/PTV methods are suitable for measuring instantaneous velocity with reasonable accuracy. Sparsity in space implies that we want to resolve scales that are small compared to the characteristic distance between measurement points. The data assimilation methods we have discussed connect successive moments in time by reproducing the dynamics determined by the Navier-Stokes equation. This allows information about the velocity field from different moments in time to be combined, increasing the accuracy and spatial resolution of possible reconstructions.
Our goal is to enhance the experimental data by increasing their spatial resolution and inferring flow variables and system parameters that were not measured directly. In particular, we aim to reconstruct the spatially dense velocity and pressure fields, establish the driving force, and unknown fluid viscosity and bottom friction coefficient, based only on velocity measurements. Knowledge of these additional quantities will allow us to better describe and characterize the properties of the considered system. The present investigation complements recent studies Eivazi, Wang, and Vinuesa (2024); Wang, Liu, and Wang (2022); Clark Di Leoni et al. (2023a); Cai, Gray, and Karniadakis (2024) by exploring a different flow configuration.
The performed analysis shows that sparse ($150$ vectors per image) but accurate velocity measurements are sufficient to reconstruct a dense ($65536$ vectors per image) velocity field with a relative root-mean-square error of about $0.2\%$ and similarly dense pressure and force fields with an accuracy several times worse for typical experimental conditions. The fluid viscosity and the bottom friction are restored with relative errors of several percent. The developed method is robust to small noise ($\leq 1\%$) in the initial velocity data and is even capable of correcting measurement errors based on the fact that the reconstructed quantities should satisfy the Navier-Stokes equation and incompressibility conditions. As the noise level in the initial data increases, the reconstruction accuracy gradually decreases. The changes in reconstruction accuracy with velocity data density are also addressed. Analysis of the energy spectra of velocity fields shows that small scales are the worst reconstructed, especially for noisy measurements.
II Governing equations
@@
We consider two-dimensional incompressible fluid flow described by the forced Navier-Stokes equation
@@
where $\bm{v}$ is 2D velocity, $p$ is the pressure, $\alpha$ is the bottom friction coefficient, and $\nu$ is the kinematic viscosity. The external force $\bm{f}$ is assumed to be stationary in time. We are interested in a moderately turbulent regime, when the nonlinear term in the Navier-Stokes equation plays a substantial role. The flow is observed in some region with open boundaries, and to study the flow we measure the velocity field. We have no information about the pressure field and the external force, except that the latter is constant in time.
It is critical to recognize that in this formulation of the problem, pressure and external force cannot be clearly separated. Indeed, by performing the Helmholtz decomposition, the external force can be represented as the sum of the gradient and solenoidal fields. The gradient contribution can be included in the definition of pressure, and then, without loss of generality, the external force in equation (1) can be considered solenoidal, i.e. $\nabla\cdot\bm{f}=0$. But even taking this condition into account, the division into pressure and external force is not unambiguous. These quantities are defined up to a transformation $p(\bm{r},t)\to p(\bm{r},t)+h(\bm{r})+c(t)$, $f(\bm{r})\to f(\bm{r})+\nabla h(\bm{r})$, where $h(\bm{r})$ is a harmonic function $\nabla^{2}h(\bm{r})=0$ in the observation zone and $c(t)$ is an arbitrary function that does not depend on spatial coordinates. An unambiguous definition of $h(\bm{r})$ and $c(t)$ is impossible without knowledge of the boundary conditions for pressure and external force at the edges of the observation region. Note that $h(\bm{r})$ does not depend on time, since the external force is assumed to be static.
The above should be kept in mind when determining the accuracy of field reconstructions. In the following, to characterize errors in the reconstructions of pressure and force fields, we will compare the values for their Laplacian $\nabla^{2}p$ and curl $\phi=\partial_{x}f_{y}-\partial_{y}f_{x}$, respectively. Knowledge of velocity dynamics allows us to determine these quantities unambiguously.
III Data collection
@@
Data for this study are obtained numerically by integrating (1) using the GeophysicalFlows.jl pseudospectral code Constantinou et al. (2021), which has recently been successfully applied to model two-dimensional turbulence Parfenyev (2022); Kolokolov, Lebedev, and Parfenyev (2024); Parfenyev (2024). The code can be executed on the GPU, resulting in high computational performance. The aliasing errors are removed with the two-third rule. The simulation domain is a doubly periodic box of size $2\pi\times 2\pi$. The velocity measurements are performed in the observation area, which is smaller and has the size of $L\times L$ with $L=\pi$, see Fig. 1a. Since the boundary conditions at the edges of the observation zone are unknown, the application of adjoint- and ensemble-variational methods for data assimilation is problematic.
The system parameters $\nu=0.01$ and $\alpha=0.1$ are chosen to be consistent with the experimental studies Xia, Shats, and Falkovich (2009). The external force has the Kolmogorov form
@@
with $f_{0}=10$ and $k_{f}=5$. We start from an arbitrary random initial condition, and after a transient process, the flow reaches a statistical steady-state. Once stationary, we saved data at $\Delta t=0.02$ intervals for $T=4$ units, covering several turn-over times of flow fluctuations. This data recording rate corresponds to video recording in a laboratory experiment at $50$ frames per second. The spatial resolution of direct numerical simulation (DNS) is $512^{2}$ for the entire computational domain and $256^{2}$ for the observation zone.
@@
Fig. 1a shows snapshot of the vorticity field $\omega=\partial_{x}v_{y}-\partial_{y}v_{x}$ in the statistical steady-state. The flow is chaotic, with no obvious imprint of an external forcing. Since there are two dissipation mechanisms – viscosity and bottom friction – the flow state can be characterized by two dimensionless parameters Mishra et al. (2015): the Reynolds number $Re=UL/\nu\approx 1.3\times 10^{3}$ and $Rh=U/(\alpha L)\approx 13$, which were computed using the root mean square velocity $U\approx 4.1$ and the size $L=\pi$ of the observation domain. The energy spectrum is presented in Fig. 1b. It indicates that fluid flow is mainly determined by modes with wave vectors $k\lesssim k_{f}$.
To imitate velocity measurements, we randomly scatter $N_{data}$ points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$ within the spatio-temporal ($256\times 256\times 200$) observation domain. We consider the case when the average distance between measurement points for a single moment in time is comparable to or larger than the forcing scale. It turns out that such sparse data are sufficient to perform reconstructions and infer system parameters with reasonable accuracy. In the experiments we rely on, the spatial resolution of the velocity measurements is usually much higher. In such a situation, it makes sense to thin out the measured velocity field by randomly selecting a relatively small number of data points, which can significantly speed up the reconstruction. The primary purpose in the above scenario is to infer the system parameters and flow variables that cannot be measured directly. If the experimental data are initially sparse in space, then the location of measurement points at close moments in time may have some spatial correlation that is not taken into account in our measurement model. The impact of such correlations on reconstruction accuracy requires further study and is beyond the scope of the present work.
The measured velocity $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$ may differ from the DNS data $\bm{v}(\bm{r}_{n}^{d},t_{n}^{d})$ due to measurement noise. We model this distortion as
$\bm{v}^{d}=\begin{pmatrix}1+\eta_{1}&0\\ 0&1+\eta_{2}\end{pmatrix}\bm{v},$ (3)
where $\eta_{i}\sim\mathcal{N}(0,\varepsilon^{2})$ are independent normally distributed random variables and the standard deviation $\varepsilon$ controls the noise level. Noise-free data correspond to $\varepsilon=0$. Note that this simple method to noise modeling accounts only for mistakes in estimating the velocity amplitude and ignores errors in wrongly referencing the velocity vector in space.
In most examples below, we set $N_{data}=3\times 10^{4}$ unless another value is explicitly specified. On average, this corresponds to $150$ measurement points per snapshot compared to $65536$ points at full resolution. The gathered data $\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})$ are used for subsequent reconstruction of the flow state and system parameters with the PINN method.
IV Physics-Informed Neural Network
@@
In this section, we describe in detail the PINN method used in our study. We utilize two independent neural networks (NNs) that are trained simultaneously, see Fig. 2a. The first NN approximates the velocity and pressure at a particular position and time. It has three input neurons and three output neurons, $(\hat{v}_{x},\hat{v}_{y},\hat{p})=\mathcal{F}_{\vec{\theta}_{1}}(x,y,t)$. The second NN is intended to predict stationary force. As a result, it contains only two input neurons and two output neurons, $(\hat{f}_{x},\hat{f}_{y})=\mathcal{G}_{\vec{\theta}_{2}}(x,y)$. The parameters $\vec{\theta}_{1}$ and $\vec{\theta}_{2}$ denote the trainable variables for NNs. In addition, we will assume that the bottom friction coefficient and the fluid kinematic viscosity are unknown to us. Since these parameters are positive, it is convenient to represent them in the form $\hat{\alpha}=e^{\lambda_{1}}$ and $\hat{\nu}=e^{\lambda_{2}}$. The scalar parameters $\lambda_{1}$ and $\lambda_{2}$ will be adjusted during PINN training along with $\vec{\theta}_{1}$ and $\vec{\theta}_{2}$.
@@
Figure 3: Comparison of DNS data (ground truth) and PINN predictions for a model trained on noise-free velocity measurements. The number of data points and collocation points are $N_{data}=3\times 10^{4}$ and $N_{eq}=9\times 10^{4}$, respectively.
@@
In general, increasing the size of NNs can reduce the prediction error, when the NNs are given enough data. Here we set $N_{l}=7$ and $N_{n}=250$ for $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$, and $N_{l}=5$ and $N_{n}=30$ for $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$. Note that $\mathcal{G}_{\vec{\theta}_{2}}(\bm{r})$ is much simpler than $\mathcal{F}_{\vec{\theta}_{1}}(\bm{r},t)$ because it is not time dependent and the external forcing usually does not have a complex spatial dependency.
Figure 4: Dependence of relative root-mean-square errors of PINN reconstructions on the number $N_{data}$ of data points used for training. Dashed lines are added for reference and their slopes are indicated. The shaded areas show the standard deviations that were estimated by training five independent PINN models for each value of $N_{data}$.
@@
$\mathcal{L}_{data}=\dfrac{1}{N_{data}U^{2}}\sum_{n=1}^{N_{data}}|\hat{\bm{v}}(\bm{r}_{n}^{d},t_{n}^{d})-\bm{v}^{d}(\bm{r}_{n}^{d},t_{n}^{d})|^{2},$ (5)
@@
$\bm{e}_{1}=\left(\partial_{t}\hat{\bm{v}}+(\hat{\bm{v}}\cdot\nabla)\hat{\bm{v}}+\nabla\hat{p}+e^{\lambda_{1}}\hat{\bm{v}}-e^{\lambda_{2}}\nabla^{2}\hat{\bm{v}}-\hat{\bm{f}}\right)L/U^{2},\;$
$e_{2}=\left(\partial_{x}\hat{v}_{x}+\partial_{y}\hat{v}_{y}\right)L/U,$
$e_{3}=\left(\partial_{x}\hat{f}_{x}+\partial_{y}\hat{f}_{y}\right)L^{2}/U^{2}.$
Computing the right-hand sides involves differentiating the outputs of NNs with respect to input neurons using automatic differentiation Baydin et al. (2018). The number and location of collocation points are not limited by measurements and can be chosen arbitrarily. We set $N_{eq}=9\times 10^{4}$, with part of these points corresponding to data points $\{\bm{r}_{n}^{d},t_{n}^{d}\}_{n=1}^{N_{data}}$, and the rest are generated randomly. Note that instead of requiring the incompressibility conditions to be satisfied, the NNs can be modified to predict the relevant stream functions Raissi, Perdikaris, and Karniadakis (2019); Kag, Seshasayanan, and Gopinath (2022). However, in this case, calculating the terms in the Navier-Stokes equation involves computation of higher-order derivatives, which significantly slows down the training process.
@@
Minimization of the loss function (7) is performed by a gradient-based optimizer that adjusts network $\{\vec{\theta}_{1},\vec{\theta}_{2}\}$ and system $\{\lambda_{1},\lambda_{2}\}$ parameters. First, we train the model for $10^{5}$ epochs using Adam optimizer Kingma and Ba (2014) with learning rate of $\mu=10^{-3}$. An epoch refers to one complete pass of the training dataset. To reduce computation time, we have implemented a mixed precision technique Micikevicius et al. (2017). Then we use L-BFGS-B optimizer Byrd et al. (1995) until the tolerance reaches machine precision. The training is carried out on a single NVIDIA A100 GPU. The code is written using the PyTorch library Paszke et al. (2019) and is openly available at [rep, ].
V Results